Identify the solution(s) of the system of equations, if any.
3x-2y=6
5x=5-25y
Possible answers :(28/13, 3/13) (-3/17, 32/17) (32/17, -3/17)(32/9, 7/3)

Answer by richwmiller(17219)

(32/17, -3/17)
x = 32/17, y = -3/17